A 3.1 sample of an unknown organic gas molecules composed of carbon, hydrogen and oxygen undergoes complete combustion to produce 4.4g CO2 and 2.7g H20. Calculate the empirical and molecular formular of the gas molecule if the molar mass is 62g/mol


1
Expert's answer
2021-04-19T05:14:36-0400

mass of C in CO2 = 12/44 × 4.4 = 1.2g

mass of H in H2O = 2/18 × 2.7 = 0.3g

mass of O in the organic compound = 3.1g - (1.2 + 0.3)g = 1.6g


C = 1.2/12 = 0.1/0.1 = 1

H = 0.3/1 = 0.3/0.1 = 3

O = 1.6/16 = 0.1/0.1 = 1


"\therefore" the empirical formula of the compound is CH3O


(CH3O)x = 62

(12 + (3×1) + 16)x = 62

x(31) = 62

x = 2

Molecular formula = (CH3O)2 = C2H6O2


"\therefore" the molecular formula of the compound is C2H6O2